OpenStudy (anonymous):

Help? Taylor has a credit card with an interest rate of 24.5% compounded monthly that she must pay off in seven years. If her balance on the credit card is $2,345, what must her monthly payment be? $27.92 $58.61 $222.31 $732.52

OpenStudy (anonymous):

m1 = B(1+r/12) - P m2 = B(1+r/12)^2 - P(1+r/12) - P m3 = B(1+r/12)^3 - P(1+r/12)^2 - P(1+r/12) - P the balance looks to be an exponential deal, while the Payments follow a geometric series

OpenStudy (anonymous):

Go on?

OpenStudy (anonymous):

since there are 7*12 periods to account for